2) For the function y = x2 - 4x - 5, perform the following tasks:
a) Put the function in the form y = a(x - h)2 + k.
Answer:
Show work in this space



b) What is the equation for the line of symmetry for the graph of this function?
Answer:



c) Graph the function using the equation in part a. Explain why it is not necessary to plot points to graph when using y = a (x – h)2 + k.
Show graph here.


Explanation of graphing.




d) In your own words, describe how this graph compares to the graph of y = x2?
Answer:

Answer by venugopalramana(3286)

SEE THIS EXAMPLE AND TRY.IF STILL IN DIFFICULTY PLEASE COME BACK
-------------------------------
2) For the function y = x2 - 6x + 8, perform the
following tasks:
a) Put the function in the form y = a(x - h)2 + k.
Answer:
Show work in this space.


b) What is the line of symmetry?
Answer:


c) Graph the function using the equation in part a.
Explain why it is not necessary to plot points to
graph when using y = a (x - h) 2 + k.
Show graph here.
Explanation of graphing.
d) In your own words, describe how this graph compares
the graph of y = x2?
Answer:
1 solutions
Answer 22402 by venugopalramana(1834)on
2006-05-06 05:49:34 
Function y = x2 - 4x - 5, perform the
following tasks:
a) Put the function in the form y = a(x - h)2 + k.
b) What is the line of symmetry?
c) Graph the function using the equation in part
a. Explain why it is not necessary to plot points to
graph when using y = a (x – h)2 + k.
Show graph here.
Explanation of graphing.
d) In your own words, describe how this graph compares
to the graph of y = x2?
1 solutions
Answer 13849 by venugopalramana(1619)on
2006-01-28 11:20:53 
Y=X^2-4X-5={(X-2)^2-4-5}=(X-2)^2-9
COMPARING WITH THE GIVEN EQN .
y=a(x-h)^2+k,WE INFER THAT
A=1,H=2 AND K=-9....THE LINE OF SYMMETRY IS X-2=0 AS
YOU WILL GET SAME VALUE OF Y WHETHER X-2=+4 SAY OR
-4...NAMELY,Y=7.
COMPARISON WITH Y=X^2 IS SHOWN BELOW
graph( 600, 600, -10, 10, -10, 10, x^2-4x-5, x^2)
YOU CAN SEE THAT LINE OF SYMMETRY IS X=0 HERE.
ALSO THE MINIMUM VALUE OR VERTEX AT 0,0 IN CASE OF
Y=X^2,WHERE AS IT WAS AT (2,-9)
FOR THE GIVEN EQUATION